Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

8 red card number 1-8 put in one box, 8 blue card number 1

Expert replies
by Max@Math Revolution » Fri Jun 10, 2016 10:14 pm
8 red card number 1-8 put in one box, 8 blue card number 1-8,put in another box, take one card from each of the box,what is the probability that the number match each other?
A. 1/2
B. 1/4
C. 1/8
D. 1/16
E. 1/64

*An answer will be posted in 2 days.
Join the discussion
Source: — Problem Solving |

by Gmatasap » Sat Jun 11, 2016 7:59 am
The probability of getting any of the number 1-8 card is 1. Since any specific number has not be provided, any number will do.
The probability of getting the same number in the draw of blue card is 1/8

The required probability is
[spoiler]1*1/8=1/8[/spoiler]

Actually modifying the question, can you help with 2 cards were drawn from each box, what is the probability that we get at least one pair of numbers?
Join the discussion

by Max@Math Revolution » Mon Jun 13, 2016 8:52 am
From (1,1),(2,2),(3,3),(4,4),(5,5),(6,6),(7,7),(8,8), the probability of the number match each other is 8. Hence, 8C1/(8C1)(8C1)=1/8. So, the correct answer is C.
Join the discussion

by 800_or_bust » Mon Jun 13, 2016 9:23 am
Max@Math Revolution wrote:8 red card number 1-8 put in one box, 8 blue card number 1-8,put in another box, take one card from each of the box,what is the probability that the number match each other?
A. 1/2
B. 1/4
C. 1/8
D. 1/16
E. 1/64

*An answer will be posted in 2 days.
The probability of selecting any card from box A is 8/8 = 1. The probability of selecting the same card from Box B is 1/8. Thus, the probability of selecting two matching cards is just 1/8.
800 or bust!
Join the discussion

by 800_or_bust » Mon Jun 13, 2016 9:24 am
Gmatasap wrote:The probability of getting any of the number 1-8 card is 1. Since any specific number has not be provided, any number will do.
The probability of getting the same number in the draw of blue card is 1/8

The required probability is
[spoiler]1*1/8=1/8[/spoiler]

Actually modifying the question, can you help with 2 cards were drawn from each box, what is the probability that we get at least one pair of numbers?
Wouldn't it just be 1/56 (assuming no replacement) or 1/64 (with replacement) using the same logic?

Edit: Nevermind, I misread your question. Thought you meant probability that both were the same. Let me think about that.

Edit 2: Hmmm, this is tricky. If the cards are replaced each time, I believe it would just be 15/64. The probability of not getting a match either time would be 7/8 x 7/8 = 49/64. So, the probability of at least one match would be 1 - 49/64 = 15/64. Without replacement, it gets tough. The first draw from Box B could have removed the card that was drawn from Box A on the second draw, in which case the probability of a match is 0. Otherwise, the probability would of a match would be 1/7. I'm guessing you would need to determine the relative probabilities of both of these cases.
800 or bust!
Join the discussion

by 800_or_bust » Mon Jun 13, 2016 9:41 am
800_or_bust wrote:
Gmatasap wrote:The probability of getting any of the number 1-8 card is 1. Since any specific number has not be provided, any number will do.
The probability of getting the same number in the draw of blue card is 1/8

The required probability is
[spoiler]1*1/8=1/8[/spoiler]

Actually modifying the question, can you help with 2 cards were drawn from each box, what is the probability that we get at least one pair of numbers?
Wouldn't it just be 1/56 (assuming no replacement) or 1/64 (with replacement) using the same logic?

Edit: Nevermind, I misread your question. Thought you meant probability that both were the same. Let me think about that.

Edit 2: Hmmm, this is tricky. If the cards are replaced each time, I believe it would just be 15/64. The probability of not getting a match either time would be 7/8 x 7/8 = 49/64. So, the probability of at least one match would be 1 - 49/64 = 15/64. Without replacement, it gets tough. The first draw from Box B could have removed the card that was drawn from Box A on the second draw, in which case the probability of a match is 0. Otherwise, the probability would of a match would be 1/7. I'm guessing you would need to determine the relative probabilities of both of these cases.
My best guess at it...

The probability that the second card drawn from Box A matches the first card drawn from Box B equals the probability that the first card drawn from Box A did not match the first card drawn from Box B, times the probability that the second card drawn from Box A does match the first card drawn from Box B given that the first card drawn from Box A did not match the first card drawn from Box B. Thus, the probability of this would be 7/8 x 1/7 = 1/8. And the probability that the second card drawn from Box A does not match the first card drawn from Box B would be 1 - 1/8 = 7/8.

Thus, the overall probability that the second cards match would be (1/8)(0) + (7/8)(1/7) = 1/8. And the probability that at least one of the two cards drawn match would again be 15/64.
800 or bust!
Join the discussion